OpenStudy (anonymous):

1 - x --- x + 1 --------- 2 - x - 1 ----- x

OpenStudy (anonymous):

\[\frac{1-\frac{x}{x+1}}{2-\frac{x-1}{x}}\]

OpenStudy (anonymous):

multiply top and bottom by \[x(x+1)\] to clear the factions

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

so numerator will be \[(1-\frac{x}{x+1})\times x(x+1)\] \[=x(x+1)-x^2\] denominator will be \[(2-\frac{x-1}{x})\times x(x+1)\] \[2x(x+1)-(x-1)(x+1)\] giving \[\frac{x(x+1)-x^2}{2x(x+1)-(x-1)(x+1)}\]

OpenStudy (anonymous):

then multiply out and combine like terms in numerator and denominator

OpenStudy (anonymous):

numerator is \[x\] denominator is \[2x^2+2x-(x^2-1)=x^2+2x+1\] if my algebra is correct.

myininaya (myininaya):

looks good

OpenStudy (anonymous):

why thank you!

myininaya (myininaya):

dont let it go to your head